ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 166, 2008

Boeckx, E. and Jong Taek Cho Pseudo-Hermitian symmetries  125

Cho, Jong Taek See E. Boeckx Chong, C. T., A. Nies and Liang Yu Lowness of higher random-

ness notions

39

D’Este, G. On partial tilting modules and bounded complexes  17
Eick, B. Schur multiplicators of finite p-groups with fixed coclass  157

Eick, B. Schur multiplicators of infinite pro-p-groups with finite coclass

147

Fernández-Alcober, G. A., J. González-Sánchez and A. Jaikin-Zapirain Omega subgroups of pro-p groups

393

Frieze, A. and M. Krivelevich On two Hamilton cycle problems in random graphs

221

González-Sánchez, J. See G. A. Fernández-Alcober Hamid, S. M., C. Onica and C. Pearcy Quasidiagonality and the

hyperinvariant subspace problem

285

Jaikin-Zapirain, A. See G. A. Fernández-Alcober Janko, Z. Finite 2-groups with exactly one nonmetacyclic maximal

subgroup

313

Juhász, I., L. Soukup and Z. Szentmiklóssy Resolvability and monotone normality

1

Krivelevich, M. See A. Frieze Li, Haizhong, Hui Ma and Linlin Su Lagrangian spheres in the

2-dimensional complex space forms

113

Ma, Hui See Haizhong Li Myers, A. N. and H. S. Wilf Left-to-right maxima in words and

multiset permutations

167

Nies, A. See C. T. Chong Onica, C. See S. M. Hamid

Paran, E. Algebraic patching over complete domains  185
Parlier, H. Fixed point free involutions on Riemann surfaces  297

Pearcy, C. See S. M. Hamid

Rosen, E. Prolongations in differential algebra  239

Rosendal, C. Automatic continuity in homeomorphism groups of compact 2-manifolds

349

Schein, M. M. Weights in Serre’s conjecture for Hilbert modular forms: the ramified case

369

Shelah, S. Vive la différence III  61

Soukup, L. See I. Juhász Su, Linlin See Haizhong Li Szentmiklóssy, Z. See I. Juhász Venkataramana, T. N. Virtual Betti numbers of compact locally

symmetric spaces

235

Wilf, H. S. See A. N. Myers Wilkens, B. Finite p-groups not characteristic in any p-group in

which they are properly contained

97

Yu, Liang See C. T. Chong